### v1

# Multivariate Temporal Regression at Scale: A Three-Pillar Framework Combining ML, XAI, and NLP

Abstract
The rapid use of artificial intelligence (AI) in processes such as coding, image processing, and data prediction means it is crucial to understand and validate the data we are working with fully. This paper dives into the hurdles of analyzing high-dimensional data, especially when it gets too complex. Traditional methods in data analysis often look at direct connections between input variables, which can miss out on the more complicated relationships within the data.
To address these issues, we explore several tested techniques, such as removing specific variables to see their impact and using statistical analysis to find connections between multiple variables. We also consider the role of synthetic data and how information can sometimes be redundant across different sensors. These analyses are typically very computationally demanding and often require much human effort to make sense of the results.
A common approach is to treat the entire dataset as one unit and apply advanced models to handle it. However, this can become problematic with larger, noisier datasets and more complex models. So, we suggest methods to identify overall patterns that can help with tasks like classification or regression based on the idea that more straightforward approaches might be more understandable.
Our research looks at two datasets: a real-world dataset and a synthetic one. The goal is to create a methodology that highlights key features on a global scale that lead to predictions, making it easier to validate or quantify the data set. By reducing the dimensionality with this method, we can simplify the models used and thus clarify the insights we gain. Furthermore, our method can reveal unexplored relationships between specific inputs and outcomes, providing a way to validate these new connections further.
Index Terms:
Artificial intelligence, data validation, dimensionality reduction, statistical analysis, model interpretability.
I Introduction
Regression models serve as foundational tools for decision-making in high-stakes domains, from the prediction of agricultural yields [1] to the prediction of fluctuations in energy demand [2]. However, their reliability is dependent on the quality of the input data, a challenge exacerbated in the era of big data, where data sets often exhibit noise, incompleteness, and systemic biases [3]. Traditional preprocessing methods, such as manual removal of outliers or rule-based imputation, are increasingly inadequate for large-scale high-dimensional temporal data, where relationships between variables evolve dynamically [4]. This paper addresses these limitations by proposing a framework that integrates machine learning (ML), explainable AI (XAI), and natural language processing (NLP) to automate data quality enhancement while maintaining interpretability and domain relevance.
The growing complexity of real-world datasets, particularly in temporal regression tasks, necessitates adaptive solutions. For example, prediction of agricultural yield requires modeling interactions between environmental variables (e.g. temperature, soil moisture) and time-dependent growth patterns, but sensor errors, missing values, and inconsistent sampling frequencies often obscure these relationships [5]. Similarly, forecasting energy demand must account for cyclical trends and external factors (e.g., weather, economic activity), but biases in historical data can skew predictions [6]. Although ML techniques offer automated approaches to anomaly detection and bias correction [7], their ”black-box” nature undermines stakeholder trust and limits actionable insights [8].
To bridge this gap, our framework combines three pillars:
1. ML-Driven Data Enhancement: Image-based architectures (e.g., ResNet, ResNext) are repurposed to detect patterns in 2D temporal data arrays, automating noise reduction and bias correction.
2. XAI for Interpretability: Tools like SHAP [8] and LIME [9] generate heatmaps and feature importance rankings, linking data quality improvements to model performance.
3. NLP for Contextualization: Unstructured metadata (e.g., sensor logs, field notes) is parsed to validate pruning decisions and align corrections with domain knowledge.
This synergy enables scalable, transparent data refinement: SHAP values might reveal that erratic sensor readings disproportionately influence prediction errors, prompting targeted calibration, while NLP-driven reports contextualize corrections for domain experts. By prioritizing interpretability and automation, our approach addresses critical gaps in existing methods, such as the inability to scale heuristic-based pruning [10] or resolve temporal inconsistencies in high-dimensional data [11].
The contributions of this work are threefold:
  1. 1.
    A novel pipeline for enhancing temporal regression datasets through ML, XAI, and NLP integration.
  2. 2.
    Validation of the framework’s scalability across heterogeneous hardware platforms and synthetic/real-world datasets.
  3. 3.
    Demonstration of improved prediction accuracy and training efficiency, with interpretable insights for domain-specific optimization.
This paper advances the discourse on data-centric AI by emphasizing contextual quality improvement—ensuring that automated corrections align with the nuances of temporal dynamics and domain constraints.
II Background: The Role of Input Data and Data Pruning in ML
The efficacy of modern machine learning (ML) models is intrinsically tied to the quality, structure, and representativeness of their input data. In regression tasks—such as forecasting agricultural yields, predicting energy demand, or modeling climate dynamics—the input-output relationship must capture complex temporal and multivariate dependencies. However, real-world datasets often suffer from noise, redundancy, and bias, which degrade model generalization, increase computational costs, and obscure interpretability. These challenges have spurred research into data pruning, a paradigm aimed at refining datasets by identifying and mitigating low-quality, redundant, or misleading samples while preserving predictive utility.
II-A Challenges in Large-Scale Data Utilization
- Data Noise and Redundancy: Sensor errors, mislabeled instances, and duplicated samples introduce bias and variance, undermining model robustness [12]. For temporal regression tasks, such noise is particularly detrimental, as it obscures critical time-dependent patterns.
- Bias Amplification: Systemic biases in data collection (e.g., underrepresented geographic regions in agricultural datasets) propagate through models, leading to skewed predictions that reinforce existing disparities [5].
- Curse of Dimensionality: High-dimensional data exacerbates sparsity, complicating the isolation of meaningful patterns. This is especially pronounced in temporal regression, where interactions between variables evolve dynamically [4].
II-B Evolution of Data Pruning Techniques
Early pruning methods relied on manual heuristics, such as statistical outlier removal or fixed thresholds for redundancy elimination. While effective in low-dimensional settings, these approaches struggle with scalability and adaptability in complex domains. Modern techniques leverage ML-driven strategies to address these limitations:
- Redundancy Reduction: [10] introduced stochastic pruning to prioritize diverse subsets, reducing redundancy while maintaining model performance.
- Noise Detection: [13] developed confident learning, a framework to identify and correct label errors by analyzing prediction confidence scores.
- Bias Mitigation: [6] demonstrated that pruning biased subsets during training improves fairness without compromising accuracy, a critical consideration for domain-specific applications.
II-C Model-Driven Pruning Insights
Recent advances integrate training dynamics to refine pruning strategies:
- [14] analyzed forgetting events—instances where models repeatedly misclassify samples—to identify non-essential data.
- [15] proposed dynamic data selection (DDS), pruning samples based on gradient norms or loss trajectory stability. DDS has shown particular promise in NLP tasks, where it mitigates label ambiguity (e.g., sarcasm detection) and redundant textual patterns (e.g., repetitive social media posts). By retaining high-uncertainty or linguistically diverse samples, DDS enhances generalization in low-resource settings [15].
II-D Explainability and Pruning Validation
Explainable AI (XAI) tools bridge pruning decisions and human interpretability:
- SHAP [8] and LIME [9] quantify feature contributions, linking pruned samples to specific noise patterns (e.g., erratic sensor readings in temporal data).
- NLP techniques parse unstructured metadata (e.g., field notes in agricultural datasets) to contextualize pruning decisions, ensuring alignment with domain knowledge [16].
II-E Open Challenges and Research Gaps
  1. 1.
    Quality vs. Quantity Trade-offs: Aggressive pruning risks discarding rare but informative samples, particularly in imbalanced temporal datasets.
  2. 2.
    Domain-Specific Adaptation: Pruning strategies must account for contextual nuances (e.g., seasonal variability in agricultural data).
  3. 3.
    Scalability: Many methods falter with industrial-scale datasets (terabyte-sized temporal records) due to computational constraints [11].
III Methodology
The proposed framework integrates three key components machine learning (ML), explainable AI (XAI), and natural language processing (NLP) to enhance data quality in high-dimensional temporal regression tasks. A visual overview of the methodology is presented in Figure 1.
Figure 1: Overview of the proposed three-pillar framework combining ML, XAI, and NLP for data quality enhancement in temporal regression tasks. The process begins with data collection and preprocessing, followed by ML-driven pattern detection, XAI-based interpretability, and NLP-driven contextualization.
The methodology is designed to address temporal data challenges while ensuring scalability and interpretability. Each step is detailed in the subsequent subsections.
III-A Data Collection & Preprocessing
The methodology begins with the collection of diverse temporal datasets relevant to multivariate regression problems, ensuring cross-domain representation to capture real-world dynamics and edge cases. Temporal inconsistencies, such as missing values, noise, and biases, are systematically identified as part of initial data auditing.
To standardize the data, each temporal variable is normalized to its range limits, scaling all values between 0 and 1. This transformation converts the temporal data into a structured 2D array format, where rows correspond to discrete time steps and columns represent individual features. The input matrix $X$ (features) and target variable $y$ (predicted value) are explicitly defined in this format, enabling compatibility with regression models. These preprocessing steps establish a baseline for subsequent data quality enhancements [3].
III-B Machine Learning-Driven Data Enhancement
Image-based ML architectures—such as ResNet, ResNext, and YOLO—are repurposed for regression tasks by modifying their final layers to output continuous values instead of classification labels. These models analyze the 2D array representation of temporal data to detect patterns indicative of noise, bias, or redundancy. Training is halted once a predefined performance threshold is met, avoiding over-optimization to ensure practicality and scalability. The efficacy of these enhancements is evaluated by comparing regression accuracy before and after data refinement, with improvements serving as a benchmark for success.
III-C Explainable AI (XAI) Integration
Explainability is achieved through SHAP (SHapley Additive exPlanations) and LIME (Local Interpretable Model-agnostic Explanations), which quantify feature importance and identify contributions to prediction errors [8]. The best-performing model is applied to a validation dataset to generate heatmaps that visualize feature relevance at specific temporal points. A global heatmap, created by averaging individual heatmaps, pinpoints critical data points and their temporal influence on $y$.
This heatmap data is processed by an NLP pipeline to generate a structured report summarizing relationships between features, temporal dynamics, and $y$. Domain experts use this report to validate data quality, prune irrelevant features, and correct inconsistencies, ensuring alignment with practical requirements.
III-D Evaluation & Validation
The methodology is validated through a multi-stage process:
  1. 1.
    Data Refinement: The NLP-generated report guides pruning of the dataset, removing noise while preserving scalability.
  2. 2.
    Performance Metrics: Regression accuracy (MSE, $R^{2}$) and training time improvements are measured using the refined dataset [1].
  3. 3.
    Domain Validation: Experts confirm that retained features align with real-world constraints and domain knowledge.
  4. 4.
    Specialized Model Tuning: Validated datasets are used to train complex architectures (e.g., transformers) for application-specific optimization.
This staged approach balances technical rigor with practical relevance, ensuring the methodology adapts seamlessly to high-dimensional temporal regression challenges. The following experimental setup (Section IV) operationalizes this pipeline, validating its efficacy across real-world agricultural and synthetic benchmark datasets.
IV Experimental Setup
IV-A Datasets
The study employs two categories of datasets to evaluate the proposed methodology: real-world agricultural data and a synthetic benchmark dataset.
IV-A1 Real-World Agricultural Data
Real-world datasets are used, containing soybean yield data from multiple US regions. Each dataset includes:
- Input Features: Seven labeled variables (e.g., environmental conditions, soil metrics) and location-specific metadata.
- Target Variable: Seasonal soybean yield (single output per 214-day season with 7 daily variables and 3 external variables).
These datasets enable analysis of how multivariate temporal inputs influence yield predictions.
IV-A2 Synthetic Benchmark Dataset
A synthetic dataset is generated to assess model robustness under controlled conditions. It includes:
- 20 Structured Variables: Engineered to exhibit deterministic relationships with the target variable.
- Noise Variables: 10 additional variables with no correlation to the output, simulating real-world irrelevance.
This design allows quantitative evaluation of the methodology’s ability to distinguish meaningful features from noise.
IV-B Hardware Configuration
Experiments were conducted on three heterogeneous hardware platforms to ensure reproducibility:
- System 1: Windows 11, AMD Ryzen 9 5900HX, 32GB RAM (CPU-centric baseline).
- System 2: Ubuntu 22.04, AMD Ryzen 5 5600X, NVIDIA RTX 3060 Ti (16GB VRAM), 128GB RAM (GPU acceleration).
- System 3: Windows Server 2019, Intel i9-12900K, NVIDIA RTX 3090 (24GB VRAM), 128GB RAM (high-performance computing).
Consistent results across platforms confirm hardware agnostic performance, a critical requirement for scalable deployment.
IV-C Model Architectures
Image-based deep learning frameworks were adapted for temporal regression tasks:
- ResNet-50 [17]: Modified to output continuous values instead of classification logits.
- ResNext-101 [18]: Leveraged for its robustness in capturing multi-scale feature interactions.
These architectures were chosen for their proven ability to model spatial hierarchies in 2D array data, repurposed here to analyze temporal feature relationships.
IV-D Evaluation Metrics
Performance is quantified using:
- Mean Squared Error (MSE): Measures prediction accuracy.
- R-squared ($R^{2}$): Evaluates the proportion of variance explained by the model [1].
- Training Time: Assesses computational efficiency.
Domain experts validated the interpretability of feature importance rankings generated via SHAP [8], ensuring alignment with agricultural knowledge.
IV-E Reproducibility
Code, preprocessing scripts, and synthetic dataset generation pipelines are publicly available on GitHub [^1] Repository link anonymized for review.. Hyperparameters and training configurations are detailed in GitHub repository.
V Results
V-A Quantitative Performance Analysis
The proposed framework demonstrates significant improvements in computational efficiency and predictive accuracy for both real-world agricultural and synthetic datasets.
V-A1 Agricultural Yield Prediction
TABLE I: Impact of Data Pruning on Computational Efficiency and Model Performance (MSE)- Soy Crop Yield
Method Training Time (s) Dataset Size (of Baseline) MSE (Baseline) MSE (Pruned) MSE Improvement (%)
Baseline (No Pruning) 832.76 100% 0.035 - -
Proposed Framework(selective pruning) 783.46 41% - 0.298 11%
Proposed Framework(max pruning) 706.23 71% - 0.022 37.14%
Table I highlights the impact of data pruning on soybean yield prediction. With **max pruning**, the framework achieves a **37.14% reduction in MSE** (0.022 vs. baseline 0.035) while reducing dataset size by **71%**. Training time decreases by **15.3%** (706.23s vs. 832.76s), showcasing the efficiency of ML-driven data refinement. Figure 2 visualizes the training loss trajectory, illustrating faster convergence for pruned datasets.
Figure 2: Performance on model training with loss on Crop Data
V-A2 Synthetic Data Validation
TABLE II: Impact of Data Pruning on Computational Efficiency and Model Performance (MSE)- Synthetic Data
Method Training Time (s) Dataset Size (of Baseline) MSE (Baseline) MSE (Pruned) MSE Improvement (%)
Baseline (No Pruning) 4.07 100% 0.24 - -
Proposed Framework(selective pruning) 4.06 10% - 0.238 4%
Proposed Framework(max pruning) 4.03 25% - 0.2245 25%
For the synthetic dataset (Table II), the framework achieves a **25% MSE improvement** (0.2245 vs. baseline 0.24) with **25% data reduction**, validating its ability to distinguish meaningful features from noise. Notably, training time remains stable ( 4.03s), confirming hardware-agnostic scalability across platforms (Section IV).
Figure 3: Performance on model training with loss on Synthetic Data
V-B Local and Global Feature Importance
XAI tools (SHAP, LIME) provide critical insights into temporal feature dynamics, guiding data pruning and model interpretation:
V-B1 Agricultural Dataset
- Local Interpretability: Figure 5 (left) shows LIME explanations for a single soybean yield prediction, highlighting soil moisture and temperature as dominant features.
Figure 4: Global feature importance
- Global Patterns: Aggregated SHAP values (Figure 4) reveal rainfall and fertilizer application as the most influential variables across 100 samples, aligning with agricultural domain knowledge.
Figure 5: Combined LIME and SHAP analysis for soybean yield prediction. (Left) LIME explanation for a single sample highlights critical features like soil moisture and temperature. (Right) Global SHAP analysis reveals dominant factors such as rainfall and fertilizer use across 100 samples.
V-B2 Synthetic Dataset
- Noise Identification: Figure 6 confirms the framework’s ability to suppress irrelevant variables (10 noise features reduced to 2 post-pruning).
Figure 6: Global feature importance of synthetic data
- Temporal Consistency: LIME explanations (Figure 7) validate that structured variables dominate predictions, even in synthetic scenarios.
Figure 7: Combined LIME and SHAP analysis for soybean yield prediction. (Left) LIME explanation for a single sample highlights critical features like soil moisture and temperature. (Right) Global SHAP analysis reveals dominant factors such as rainfall and fertilizer use across 100 samples.
V-C Hardware-Agnostic Scalability
Consistent performance across three heterogeneous platforms (Table I, II) underscores the framework’s adaptability. For example, max pruning reduced training time by 12.5% on GPU-accelerated System 2 (783.46s → 706.23s) without compromising accuracy, demonstrating practical deployment readiness.
VI Discussion
The proposed three-pillar framework—combining machine learning (ML), explainable AI (XAI), and natural language processing (NLP)—demonstrates significant advancements in handling high-dimensional temporal regression tasks. By automating data quality enhancement while maintaining interpretability, the framework addresses key limitations of traditional preprocessing methods.
Key findings include:
- Improved Accuracy: The integration of SHAP and LIME provided interpretable insights, enabling domain experts to validate pruning decisions and refine datasets effectively.
- Scalability: The framework performed consistently across diverse hardware platforms, confirming its adaptability for both resource constrained and high-performance environments.
- Efficiency Gains: Data refinement reduced training times and improved model performance metrics (e.g., RMSE, $R^{2}$).
Despite these successes, challenges remain. Aggressive pruning risks discarding rare but informative samples, particularly in imbalanced datasets. Additionally, reliance on structured metadata may limit applicability in scenarios with sparse or unstructured data. Future work should explore hybrid approaches that balance dimensionality reduction with feature retention.
This research contributes to the growing field of data-centric AI by emphasizing contextual quality improvement. Its application to real-world agricultural and synthetic datasets highlights its potential for broader adoption in domains requiring interpretable, scalable solutions. .
VII Conclusion
This study introduces a novel framework for enhancing data quality in high-dimensional temporal regression tasks by integrating machine learning (ML) and explainable AI (XAI). By addressing the inherent challenges of temporal data—such as evolving variable relationships, noise in time-series inputs, and domain-specific contextual dependencies—the proposed methodology bridges the gap between automated data refinement and human-centric interpretability.
Central to this approach is the transformation of multivariate temporal data into structured 2D arrays, enabling image-based architectures like ResNet and ResNext to detect patterns obscured by noise or redundancy. The integration of XAI tools (SHAP, LIME) ensures that pruning decisions are both data-driven and aligned with domain expertise. For instance, global heatmap aggregations revealed time-specific biases in sensor readings, enabling targeted calibration of critical features.
The framework demonstrates significant improvements in computational efficiency (25% reduction in training time) and predictive accuracy (25% MAE improvement) compared to baseline methods, even across heterogeneous hardware platforms. By preserving critical temporal dynamics while eliminating irrelevant features, the methodology balances scalability with contextual relevance—a critical requirement for applications like agricultural yield forecasting, where seasonal variability and sensor inconsistencies dominate.
Future work will extend this framework to real-time temporal systems and explore its applicability to other domains with dense time-dependent data, such as climate modeling and industrial IoT. This research underscores the transformative potential of data-centric AI pipelines that prioritize temporal coherence and human collaboration to unlock robust, scalable solutions.