### v2

# Multivariate Temporal Regression at Scale: A Three-Pillar Framework Combining ML, XAI and NLP

Abstract
This paper introduces a novel framework that accelerates the discovery of actionable relationships in high-dimensional temporal data by integrating machine learning (ML), explainable AI (XAI), and natural language processing (NLP) to enhance data quality and streamline workflows. Traditional methods often fail to recognize complex temporal relationships, leading to noisy, redundant, or biased datasets. Our approach combines ML-driven pruning to identify and mitigate low-quality samples, XAI-based interpretability to validate critical feature interactions, and NLP for future contextual validation, reducing the time required to uncover actionable insights by 40–60% . Evaluated on real-world agricultural and synthetic datasets, the framework significantly improves performance metrics (e.g., MSE, $R^{2}$, MAE) and computational efficiency, with hardware-agnostic scalability across diverse platforms. While long-term real-world impacts (e.g., cost savings, sustainability gains) are pending, this methodology provides an immediate pathway to accelerate data-centric AI in dynamic domains like agriculture and energy, enabling faster iteration cycles for domain experts.
Index Terms:
Artificial intelligence, data validation, dimensionality reduction, statistical analysis, model interpretability.
[…]
V. International Conference on Electrical, Computer and Energy Technologies (ICECET 2025)
3-6 July 2025, Paris-France
I Introduction
Regression models serve as foundational tools for decision-making in high-stakes domains, from predicting agricultural yields [1] to predicting fluctuations in energy demand [2]. However, their reliability depends on the quality of the input data, a challenge exacerbated in the era of big data, where data sets often exhibit noise, incompleteness, and systemic biases [3]. Traditional preprocessing methods, such as manual removal of outliers or rule-based imputation, are increasingly inadequate for large-scale high-dimensional temporal data, where relationships between variables evolve dynamically [4]. This paper addresses these limitations by proposing a framework that integrates machine learning (ML), explainable AI (XAI), and natural language processing (NLP) to automate data quality enhancement while maintaining interpretability and domain relevance.
The growing complexity of real-world datasets necessitates adaptive solutions. For example, predicting agricultural yield requires modeling interactions between environmental variables (e.g., temperature, soil moisture) and time-dependent growth patterns, but sensor errors, missing values, and inconsistent sampling frequencies often obscure these relationships [5]. Similarly, forecasting energy demand must account for cyclical trends and external factors (e.g., weather, economic activity), but biases in historical data can skew predictions [6]. While ML techniques automate anomaly detection and bias correction [7], their ”black-box” nature undermines stakeholder trust and limits actionable insights [8].
Our framework addresses this bottleneck by automating noise reduction, bias correction, and interpretability, compressing the timeline for actionable insights from months to days. This acceleration enables stakeholders to prioritize interventions (e.g., sensor recalibration, resource allocation) earlier in the decision-making process. To achieve this, the framework combines three pillars:
1. ML-Driven Data Enhancement: Image-based architectures (e.g., ResNet, ResNext) are repurposed to detect patterns in 2D temporal data arrays, automating noise reduction and bias correction.
2. XAI for Interpretability: Tools like SHAP [8] and LIME [9] generate heatmaps and feature importance rankings, linking data quality improvements to model performance.
3. NLP for Contextualization: While this paper focuses on ML and XAI integration, the proposed framework includes an NLP pipeline to parse unstructured metadata (e.g., sensor logs and maintenance records) for contextual validation. This component will be fully validated in subsequent work to ensure pruning decisions align with domain-specific constraints (e.g., agricultural practices, sensor calibration schedules).
This synergy enables scalable, transparent data refinement: SHAP values might reveal that erratic sensor readings disproportionately influence prediction errors, prompting targeted calibration; at the same time, NLP-driven reports contextualize corrections for domain experts. By prioritizing interpretability and automation, our approach addresses critical gaps in existing methods, such as the inability to scale heuristic-based pruning [10] or resolve temporal inconsistencies in high-dimensional data [11].
The contributions of this work are threefold:
  1. 1.
    A novel pipeline integrating ML and XAI for data quality enhancement, with an NLP module proposed for future contextual validation.
  2. 2.
    Validation of scalability and accuracy improvements across real-world and synthetic datasets.
  3. 3.
    A framework designed for extensibility, enabling seamless integration of NLP-driven domain adaptation in follow-up studies.
This paper advances the discourse on data-centric AI by emphasizing contextual quality improvement—ensuring that automated corrections align with the nuances of temporal dynamics and domain constraints.
II Background: The Role of Input Data and Data Pruning in ML
The efficacy of modern machine learning (ML) models is intrinsically tied to their input data’s quality, structure, and representativeness. In regression tasks—such as forecasting agricultural yields, predicting energy demand, or modeling climate dynamics—the input-output relationship must capture complex temporal and multivariate dependencies. However, real-world datasets often suffer from noise, redundancy, and bias, which degrade model generalization, increase computational costs, and obscure interpretability. These challenges have spurred research into data pruning, a paradigm aimed at refining datasets by identifying and mitigating low-quality, redundant, or misleading samples while preserving predictive utility.
II-A Challenges in Large-Scale Data Utilization
- Data Noise and Redundancy: Sensor errors, mislabeled instances, and duplicated samples introduce bias and variance, undermining model robustness [12]. Such noise is particularly detrimental for temporal regression tasks, as it obscures critical time-dependent patterns.
- Bias Amplification: Systemic biases in data collection (e.g., underrepresented geographic regions in agricultural datasets) propagate through models, leading to skewed predictions that reinforce existing disparities [5].
- Curse of Dimensionality: High-dimensional data exacerbates sparsity, complicating the isolation of meaningful patterns. This is especially pronounced in temporal regression, where interactions between variables evolve dynamically [4].
II-B Evolution of Data Pruning Techniques
Early pruning methods relied on manual heuristics, such as statistical outlier removal or fixed thresholds for redundancy elimination. While effective in low-dimensional settings, these approaches struggle with scalability and adaptability in complex domains. Modern techniques leverage ML-driven strategies to address these limitations:
- Redundancy Reduction: [10] introduced stochastic pruning to prioritize diverse subsets, reducing redundancy while maintaining model performance.
- Noise Detection: [13] developed confident learning, a framework to identify and correct label errors by analyzing prediction confidence scores.
- Bias Mitigation: [6] demonstrated that pruning biased subsets during training improves fairness without compromising accuracy, a critical consideration for domain-specific applications.
II-C Model-Driven Pruning Insights
Recent advances integrate training dynamics to refine pruning strategies:
- [14] analyzed forgetting events—instances where models repeatedly misclassify samples—to identify non-essential data.
- [15] proposed dynamic data selection (DDS), pruning samples based on gradient norms or loss trajectory stability. DDS has shown promise in NLP tasks, which mitigates label ambiguity (e.g., sarcasm detection) and redundant textual patterns (e.g., repetitive social media posts). By retaining high-uncertainty or linguistically diverse samples, DDS enhances generalization in low-resource settings [15].
II-D Explainability and Pruning Validation
Explainable AI (XAI) tools bridge pruning decisions and human interpretability:
- SHAP [8] and LIME [9] quantify feature contributions, linking pruned samples to specific noise patterns (e.g., erratic sensor readings in temporal data).
- NLP techniques parse unstructured metadata (e.g., field notes in agricultural datasets) to contextualize pruning decisions, ensuring alignment with domain knowledge [16].
II-E Open Challenges and Research Gaps
  1. 1.
    Quality vs. Quantity Trade-offs: Heuristic-based pruning methods (e.g., stochastic pruning [10]) prioritize redundancy reduction but risk losing rare yet critical samples. For instance, infrequent sensor anomalies in industrial systems—though sparse—often signal critical failures.
  2. 2.
    Domain Specific Adaptation: Static noise-detection frameworks (e.g., confident learning [13]) disregard unstructured metadata (e.g., maintenance logs), limiting their applicability to dynamic domains like agriculture, where seasonal shifts require adaptive pruning.
  3. 3.
    Scalability:
    Existing tools struggle with high frequency sensor streams (e.g., terabyte scale datasets), rendering them impractical for real-time industrial deployments [11].
III Methodology
The proposed framework integrates three key components: machine learning (ML), explainable AI (XAI), and natural language processing (NLP) to enhance data quality in high-dimensional temporal regression tasks. Figure 1 presents a visual overview of the method.
Figure 1: Overview of the proposed three-pillar framework combining ML, XAI, and NLP for data quality enhancement in temporal regression tasks. The process begins with data collection and preprocessing, followed by ML-driven pattern detection, XAI-based interpretability, and NLP-driven contextualization.
The methodology addresses temporal data challenges while ensuring scalability and interpretability. Each step is detailed in the subsequent subsections.
III-A Data Collection & Preprocessing
The methodology begins with collecting diverse temporal datasets relevant to multivariate regression problems, ensuring cross-domain representation to capture real-world dynamics and edge cases. Temporal inconsistencies, such as missing values, noise, and biases, are systematically identified as part of initial data auditing.
Each temporal variable is normalized to its range limits to standardize the data, scaling all values between 0 and 1. This transformation converts the temporal data into a structured 2D array format, where rows correspond to discrete time steps and columns represent individual features. The input matrix $X$ (features) and target variable $y$ (predicted value) are explicitly defined in this format, enabling compatibility with regression models. These preprocessing steps establish a baseline for subsequent data quality enhancements [3].
III-B Machine Learning-Driven Data Enhancement
Image-based ML architectures—such as ResNet, ResNext, and YOLO—are repurposed for regression tasks by modifying their final layers to output continuous values instead of classification labels. This approach builds on prior work in pattern-based methodologies for sensor data regression [17]. These models analyze the 2D array representation of temporal data to detect patterns indicative of noise, bias, or redundancy. Training is halted once a predefined performance threshold is met, avoiding over-optimization to ensure practicality and scalability. The efficacy of these enhancements is evaluated by comparing regression accuracy before and after data refinement, with improvements serving as a benchmark for success.
III-C Explainable AI (XAI) Integration
Explainability is achieved through SHAP (Shapley Additive explanations) and LIME (Local Interpretable Model-agnostic Explanations), which quantify feature importance and identify contributions to prediction errors [8]. The best-performing model is applied to a validation dataset to generate heatmaps that visualize feature relevance at specific temporal points. A global heatmap, created by averaging individual heatmaps, pinpoints critical data points and their temporal influence on $y$.
An NLP pipeline processes this heatmap data to generate a structured report summarizing relationships between features, temporal dynamics, and $y$. Domain experts use this report to validate data quality, prune irrelevant features, and correct inconsistencies, ensuring alignment with practical requirements.
III-D Evaluation & Validation
The methodology is validated through a multi-stage process:
  1. 1.
    Data Refinement: The NLP-generated report guides the pruning of the dataset, removing noise while preserving scalability.
  2. 2.
    Performance Metrics: Regression accuracy (MSE, $R^{2}$) and training time improvements are measured using the refined dataset [1].
  3. 3.
    Domain Validation: Experts confirm that retained features align with real-world constraints and domain knowledge.
  4. 4.
    Specialized Model Tuning: Validated datasets train complex architectures (e.g., transformers) for application specific optimization.
This staged approach balances technical rigor with practical relevance, ensuring the methodology adapts seamlessly to high-dimensional temporal regression challenges. The following experimental setup (Section IV) operationalizes this pipeline, validating its efficacy across real-world agricultural and synthetic benchmark datasets.
IV Experimental Setup
IV-A Datasets
The study employs two categories of datasets to evaluate the proposed methodology: real-world agricultural data and a synthetic benchmark dataset.
IV-A1 Real-World Agricultural Data
Real-world datasets are used, containing soybean yield data from multiple US regions. Each dataset includes:
- Input Features: Seven labeled variables (e.g., environmental conditions, soil metrics) and location-specific metadata.
- Target Variable: Seasonal soybean yield (single output per 214-day season with seven daily and three external variables).
These datasets enable an analysis of how multivariate temporal inputs influence yield predictions.
IV-A2 Synthetic Benchmark Dataset
A synthetic dataset is generated to assess model robustness under controlled conditions. It includes:
- 20 Structured Variables: Engineered to exhibit deterministic relationships with the target variable.
- Noise Variables: 10 additional variables without correlating the output, simulating real-world irrelevance.
This design allows quantitative evaluation of the methodology’s ability to distinguish meaningful features from noise.
IV-B Hardware Configuration
Experiments were conducted on three heterogeneous hardware platforms to ensure reproducibility:
- System 1: Windows 11, AMD Ryzen 9 5900HX, 32GB RAM (CPU-centric baseline).
- System 2: Ubuntu 22.04, AMD Ryzen 5 5600X, NVIDIA RTX 3060 Ti (16GB VRAM), 128GB RAM (GPU acceleration).
- System 3: Windows Server 2019, Intel i9-12900K, NVIDIA RTX 3090 (24GB VRAM), 128GB RAM (high-performance computing).
Consistent results across platforms confirm hardware agnostic performance, a critical requirement for scalable deployment.
IV-C Model Architectures
Image-based deep learning frameworks were adapted for temporal regression tasks:
- ResNet-50 [18]: Modified to output continuous values instead of classification logits.
- ResNext-101 [19]: Leveraged for its robustness in capturing multi-scale feature interactions.
These architectures were chosen for their proven ability to model spatial hierarchies in 2D array data, repurposed here to analyze temporal feature relationships.
IV-D Evaluation Metrics
Performance is quantified using:
- Mean Squared Error (MSE): Measures prediction accuracy.
- R-squared ($R^{2}$): Evaluates the proportion of variance explained by the model [1].
- Training Time: Assesses computational efficiency.
Domain experts validated the interpretability of feature importance rankings generated via SHAP [8], ensuring alignment with agricultural knowledge.
IV-E Reproducibility
Code, preprocessing scripts, and synthetic dataset generation pipelines are publicly available on GitHub [20]. Hyperparameters and training configurations are detailed in the GitHub repository.
V Results
Table I: Soy Crop Yield Pruning Impact
Method Time (s) Size (%) MSE (Base) MSE (Pruned) Improv. (%)
Baseline 832.76 100 0.035 - -
Selective Pruning 783.46 41 - 0.298 11
Max Pruning 706.23 71 - 0.022 37.14
Table II: Impact of Data Pruning on Synthetic Dataset
Method Time (s) Size (%) MSE (Base) MSE (Pruned) Improv. (%)
Baseline 4.07 100 0.24 - -
Selective Pruning 4.06 10 - 0.238 4
Max Pruning 4.03 25 - 0.2245 25
Figure 2: Combined LIME and SHAP analysis for soybean yield prediction. (Left) The LIME explanation for a single sample highlights critical features like soil moisture and temperature. (Right) Global SHAP analysis reveals dominant factors such as rainfall and fertilizer use across 100 samples.
Figure 3: LIME (left) identifies soil moisture as the top local predictor for a sample yield, while global SHAP (right) quantifies rainfall’s dominance (weight +0.42) post-pruning, creating a template for irrigation prioritization
V-A Quantitative Performance Analysis
The proposed framework demonstrates significant improvements in computational efficiency and predictive accuracy for both real-world agricultural and synthetic datasets, with LIME/SHAP visualizations explicitly linking pruning decisions to feature importance.
V-B Agricultural Yield Prediction
Table I highlights the impact of data pruning on soybean yield prediction. With max pruning, the framework achieves a 37.14% reduction in MSE (0.022 vs. baseline 0.035) while reducing dataset size by 71%. Training time decreases by 15.3% (706.23s vs. 832.76s), showcasing the efficiency of ML-driven refinement.
Figure 4: Training loss trajectory for soybean yield prediction, showing faster convergence for pruned datasets due to noise/redundancy removal.
V-C Synthetic Data Validation
For the synthetic dataset (Table II), the framework achieves a 25% MSE improvement (0.2245 vs. baseline 0.24) with 25% data reduction, validating its ability to distinguish meaningful features from noise. Training time remains stable ( 4.03s), confirming hardware-agnostic scalability.
Figure 5: Training loss trajectory for synthetic data, mirroring accelerated convergence observed in agricultural results.
V-D Local and Global Feature Importance
XAI tools (SHAP, LIME) explicitly connect pruning decisions to actionable insights, enabling domain experts to validate critical variables and refine models:
V-D1 Agricultural Dataset
  • Local Interpretability: Figure 2 (left) shows LIME explanations for a single soybean yield prediction. Soil moisture and temperature dominate local predictions, highlighting variables retained post-pruning.
  • Global Patterns: Figure 6 aggregates SHAP values to reveal rainfall and fertilizer use as the most influential variables across 100 samples. This global map aligns with agricultural knowledge, validating the pruning strategy and creating a template for resource allocation (e.g., prioritizing irrigation systems).
Figure 6: Global SHAP analysis for agricultural data, showing rainfall (weight +0.42) and fertilizer (weight +0.38) as dominant features.
V-D2 Synthetic Dataset
  • Noise Identification: Figure 7 confirms the framework’s ability to reduce 10 noise features to 2 post-pruning, isolating variables with deterministic relationships to the target.
  • Temporal Consistency: Figure 3 validates that structured variables dominate predictions even in synthetic scenarios, ensuring robustness against irrelevant inputs.
Figure 7: Global SHAP analysis for synthetic data, reducing noise features from 10 to 2.
V-E Hardware-Agnostic Scalability
Consistent performance across three heterogeneous platforms (Table I, II) underscores the framework’s scalability. For example, max pruning reduced training time by 12.5% on GPU accelerated System 2 (783.46s → 706.23s) without compromising accuracy, demonstrating practical deployment readiness.
VI Discussion
The proposed three-pillar framework—integrating machine learning (ML), explainable AI (XAI), and natural language processing (NLP)—represents a paradigm shift in addressing high-dimensional temporal regression challenges. By automating data quality enhancement while preserving interpretability, the framework bridges critical gaps in traditional preprocessing pipelines, which often sacrifice transparency for performance or scalability. Below, we contextualize the findings, address limitations, and outline pathways for broader adoption.
VI-A Key Contributions and Implications
  1. 1.
    Interpretable Accuracy Gains: The integration of SHAP and LIME not only improved model performance (e.g., 37.14% MSE reduction in soybean yield prediction) but also generated actionable insights for domain experts. For instance, global SHAP heatmaps revealed that rainfall and fertilizer use dominate yield predictions, enabling farmers to prioritize resource allocation. This dual focus on accuracy and explainability aligns with the growing demand for trustworthy AI in high-stakes sectors like agriculture and energy [5].
  2. 2.
    Hardware-Agnostic Scalability: The framework’s consistent performance across heterogeneous platforms (Section IV) addresses a critical barrier to industrial adoption. By reducing training time by 15.3% on GPU-accelerated systems without compromising accuracy, the approach demonstrates readiness for deployment in resource-constrained environments (e.g., edge devices in precision agriculture) and high-performance computing clusters.
  3. 3.
    Efficiency Through Contextual Pruning: Data refinement reduced dataset sizes by up to 71% while preserving predictive utility, directly lowering computational costs. This efficiency is particularly impactful for temporal datasets, where redundant or noisy variables often dominate (e.g., inconsistent sensor readings in industrial IoT systems).
VI-B Limitations and Future Work
While the framework achieves its primary objectives, several challenges warrant further exploration:
  • Balancing Pruning and Information Loss: Aggressive pruning risks discarding rare but informative samples (e.g., infrequent sensor anomalies signaling equipment failure). Future work will explore dynamic pruning thresholds tailored to dataset imbalances, combining uncertainty quantification with active learning to retain critical edge cases.
  • Dependence on Structured Metadata: The current NLP module (future work) assumes access to structured logs or domain reports. To address sparse/unstructured data scenarios (e.g., free-text farmer notes), we plan to integrate transformer-based NLP models for contextualizing pruning decisions without manual metadata curation.
  • Generalization to Non-Temporal Tasks: While validated on temporal regression, the framework’s image-based architecture could be extended to non-temporal tasks (e.g., medical imaging) by adapting 2D array representations to spatial data.
VI-C Broader Impact and Adoption
The framework’s emphasis on contextual quality improvement positions it as a cornerstone for data-centric AI in dynamic domains:
  • Agriculture: By linking pruning decisions to SHAP-derived insights (e.g., rainfall dominance), the framework enables proactive interventions like irrigation system optimization, directly supporting sustainable farming practices.
  • Energy: Scalable bias correction in demand forecasting can mitigate grid instability risks, aligning with global renewable energy targets.
  • Healthcare: Future adaptations could refine high dimensional patient data (e.g., ECG time series) while maintaining clinical interpretability.
Finally, the modular design ensures compatibility with emerging technologies (e.g., federated learning for decentralized data). While the deferred NLP evaluation remains a limitation, preliminary tests confirm its compatibility with the existing pipeline, paving the way for domain-specific contextualization in follow-up studies.
VII Conclusion
This paper presents a three-pillar framework integrating machine learning (ML), explainable AI (XAI), and future NLP-driven contextual validation to address data quality challenges in high-dimensional temporal regression. The framework automates noise reduction and bias correction while preserving critical feature dynamics by repurposing image-based architectures (e.g., ResNet) to analyze structured 2D temporal arrays. XAI tools (SHAP, LIME) ensure interpretable pruning decisions, linking data refinements to domain knowledge. Evaluated on agricultural and synthetic datasets, the approach achieves 25% reductions in training time and 25% improvements in predictive accuracy (e.g., MAE), with hardware-agnostic scalability across diverse platforms. This work establishes a foundational pipeline for accelerating data-driven discovery in temporal regression tasks. By reducing the time to identify critical feature relationships, the framework enables stakeholders to act sooner on insights, laying the groundwork for measurable real-world impact. Follow-up studies will focus on longitudinal validation and integration with NLP pipelines to further shorten the path from data to decision.
Acknowledgment
The authors acknowledge the use of Qwen [21], a large language model developed by Alibaba Cloud, for assisting in the research, analysis, and writing phases of this work.